# Large Scale Multi-Task Bayesian Optimization with Large Language Models

### v1

@@ 3.4 Self-Augmentation
0:  Set of tasks $\mathcal{T}$, fine-tuned LLM $\pi_{\theta}$, iteration count $T$, criteria $\mathcal{C}$
0:  Fine-tuned LLM $\pi_{\theta+T}$
1:  for iteration $k=1$ to $T$ do
2:     for task $t\in\mathcal{T}$ do
3:        $X_{init}\sim\pi_{\theta+k}(t)$Generate samples from optimized LLM $\pi_{\theta+k}$}
4:        $X^{\star}_{init}\leftarrow\mathcal{C}\{x\,|\,x\in\mathcal{X}^{\star}_{init}\}$Select best samples using the criteria for further fine-tuning}
5:     end for
6:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{init}\}$Augment with top solutions}
7:     Fine-tune $\pi_{\theta+k}$ on $\mathcal{D}$ using instruction prompting
8:  end for
@@ 4 Experiments
We evaluate BOLT on two distinct problem domains, each with a large number of similar tasks. Both domains allow their problem definitions and solutions to be represented as strings. This allows BOLT to operate both in sequence space, where the LLM learns from optimization trajectories, and in latent space, where Bayesian optimization (BO) makes additional progress using LLM-sampled initializations.
4.1 Experimental Setup
@@
For the inner optimization loop, we implement a constrained version of the LOL-BO algorithm (Maus et al., 2022) using BoTorch and GPyTorch (Balandat et al., 2020; Gardner et al., 2018). Code to reproduce all results will be publicly available on GitHub after acceptance, with a link to the repository at placeholder. For query optimization, we use an acquisition batch size of 1 with a budget of 4,000 oracle calls, while for peptide design, we employ a larger acquisition batch size of 50 with a budget of 200,000 oracle calls.
@@
The outer loop uses instruction prompting (Mishra et al., 2021; Longpre et al., 2023) to guide the LLM in producing optimized sequences. Figure 1 shows the prompt used to prompt gpt-4o-mini for efficient query plans. After each optimization iteration, we augment the training set with the highest-scoring sequences from the optimization trajectory and fine-tune gpt-4o-mini on this expanded dataset. When fine-tuning the LLM for the query plan optimization task, we use OpenAI’s automatic batch size selection option. For the peptide design task, we found that using the automatic batch size option did not provide a similar boost in performance, and we use a constant batch size of $10$. For both tasks, we fine-tuned the LLM for $2$ epochs and used the default OpenAI LR multiplier hyperparameter of $1.8$. To ensure the solutions always have the correct syntax, we filter out characters that do not correspond to strings of integers or valid amino acids for the respective tasks. Appendix A provides additional details on the fine-tuning process and prompts used for the peptide task.
@@
Query optimization task. This task focuses on finding query plans (including join orderings and their operators) that minimize execution time for a given query. We take a subset of $2933$ queries from the Cardinality Estimation Benchmark introduced by Negi et al. (2021), keeping $99$ queries for validation. Following Tao et al. (2025), we allow BO over query plans by encoding join orders and operators as integer lists, which are then mapped to a 64-dimensional continuous latent space using the pre-trained query plan VAE from Tao et al. (2025). For the pretraining dataset, we randomly generate $1,169,890$ query plans generated based on the database schema, separated into 80/10/10 splits. For the initial “from scratch” runs with no LLM, we initialize with the set of 50 query plans used by BAO (Marcus et al., 2021), a default query optimizer that produces reasonable but non-optimal plans. Subsequent runs use 50 LLM-sampled query plans per query as initialization points. The “task description context” $C$ used to fine-tune the LLM for this task is the full SQL query string.
@@
Antimicrobial peptide design task. For this application, we are given a library of 1000 extinct, weakly antimicrobial seed peptides $S=\left\{s_{1},...,s_{L}\right\}$. A task in this setting is to take a particular seed peptide $s_{i}$ and make modifications to it to minimize the minimum inhibitory concentration (MIC) against A. Baumannii ATCC $19606$, measured in $\mu\text{mol}/L$. We created a library of $L=1000$ extinct peptides and held out the last $100$ as validation. We ensure edited peptides maintain a minimum $75\%$ similarity to the seed peptide, defined by $1-\frac{d(S,S^{\prime})}{\text{len}(S)}$, where $d$ is the Levenshtein distance between them. All of the validation peptides are at least $25\%$ different from any other peptide in the library. Although the seed peptides don’t achieve low MICs, the hope is that bacteria are less likely to have developed resistance to their variations as they come from extinct species (Wan et al., 2024). We assess MICs with the APEX model Wan et al. (2024) and utilize a VAE trained on 4.5 million amino acid sequences Torres et al. (2024) to map peptides into a $256$-dimensional latent space. Initial optimization uses $1000$ randomly mutated sequences with a similarity constraint of $75\%$ to the seed. Subsequent runs utilize $1000$ LLM sampled peptides. We use the seed amino acid sequence as the “task description context” $C$ for LLM fine-tuning.
@@
Baselines. We compare BOLT against two baseline approaches. First, we compare to “from scratch,” single task LOL-BO which we will refer to as STBO, which operates without prior task knowledge. Second, we compare to a common strategy for multi task BO—see, e.g., Patacchiola et al. (2020); Hakhamaneshi et al. (2022); Perrone et al. (2018), where a shared GP is trained on all tasks through a neural network feature extractor using the optimization trajectories from training tasks. This shared GP is then check-pointed and used on test tasks. Several papers have found success with variations of this approach. ABLR (Perrone et al., 2018) uses independent Bayesian linear regression heads per task on the shared feature extractor, while FSBO (Wistuba & Grabocka, 2021) uses an adaptation of DKT (Patacchiola et al., 2020). In this paper, for the final supervised model, we use the same PPGPR model as the BO inner-loop in our method as we require scalability but find this results in better performance than Bayesian linear regression or random Fourier feature models (Rahimi & Recht, 2007). Additional details can be found in Section B.1.
@@ 4.2 Optimization Results
Initializing Bayesian optimization with BOLT. To initialize BO with BOLT, we sample 50 points for query optimization tasks and 1,000 points for peptide optimization tasks. All sampling is performed with a temperature parameter of 0.7 for the query optimization task and 1.0 for the peptide design task unless otherwise specified. In Figure 2, we demonstrate that initializing BO with BOLT significantly improves optimization efficiency across both domains. On the query optimization task (left), while DKT/FSBO makes improvements over STBO, the gains appear to plateau after only 20 tasks. In contrast, BOLT successfully scales to over 1400 tasks and converges to far higher quality solutions faster. On the peptide design task (right), BOLT shows similarly strong performance, while DKT/FSBO struggles to take advantage of the data collected for separate templates. Surprisingly, we find that on both tasks once BOLT reaches a sufficient scale, it begins to generate initialization data for BO that is significantly better performing than the final results found by STBO and DKT/FSBO.
@@
BOLT as a one- and few-shot optimizer. As illustrated in Figure 3, BOLT demonstrates strong few-shot generalization capabilities, even achieving single-shot performance competitive with traditional systems.
In query optimization, all BOLT variants outperform the top BAO solution within 5 samples. Notably, BOLT-1138 and BOLT-1426 surpass PostgreSQL in a single sample, indicating their potential for rapid deployment in low latency scenarios. The performance of BOLT consistently improves with more iterations across both tasks, except at 50 samples on the query optimization task, where BOLT‐1138 slightly outperforms BOLT‐1426. This may be due to variances in LLM sample generation or training. Overall, these findings confirm our approach’s robustness and efficiency when scaling BOLT to thousands of tasks.
We further compare our few‐shot performance (Figure 3, before the x-axis break) to full BO runs (Figure 3 after the x-axis break). In both tasks, BOLT achieves few shot results comparable to the full BO runs.
Figure 3: Evaluating BOLT in the few shot setting and comparing to full optimization runs in both problem settings (Left: query plan optimization; Right: peptide design). In each plot, we show objective values accumulated across all validation tasks for various methods. Before the break in each plot, we show the few shot performance of BOLT and relevant domain specific few shot baselines (the PostgreSQL one-shot query optimizer, Bao). After the break in each plot, we show the results at the end of various full BO runs, which are often comparable to BOLT’s few shot performance.
@@ 4.3 Ablation Studies
In this section, we provide additional experiments analyzing various components and extensions of BOLT.
Improving LLM performance via self-augmentation. We investigate whether the performance improvements of self-augmentation as outlined in Algorithm 3 can improve the LLM while avoiding the computational expense of inner-loop BO. We apply the self-augmentation process to BOLT-1138 and BOLT-1426, generating 10 samples from each across all $2,933$ training tasks, keeping only queries that outperform the best query plan from BAO’s solutions. We then use these two dataset as additional fine-tuning to create self-augmented versions of the models.
Table 2 shows that self-augmentation yields substantial improvements even without additional tasks optimized by BO. Both self-augmented models converge to a similar performance level, achieving a summed runtime of about 62 seconds across the 100 validation queries. This convergence suggests a natural performance plateau after training on either 1,500 tasks (BOLT-1426) or 1,100 tasks plus self-generated samples (BOLT-1138+SA). The consistency of this plateau across different training approaches further demonstrates BOLT’s robustness when scaling to large task sets. This self-augmentation experiment indicates that once the LLM has been fine-tuned to sufficient performance, it can generate additional fine-tuning data, reducing the number of actual BO optimization runs required. Additionally, our framework scales well with more training tasks without performance loss.
Ablating the impact of temperature on performance. In all prior experiments, we used a constant sampling temperature of 0.7 for response generation. We demonstrate that greedy decoding (i.e., setting the temperature to 0) improves one-shot performance. As shown in Figure 4, this strategy generates a larger set of queries that surpasses the best among 50 BAO-generated queries by a wider margin.
@@
Table 1: Comparing LLMs fine tuned with (Left) data from 1138 tasks, (Right) data from 1426 tasks, and (Middle) data from 1138 tasks, but including the extra data from BOLT-1426, and removing data from older tasks. See ablation study #3 for details.
Ablating the impact of data quality on training. We perform an ablation to assess the importance of using ”better” versus ”more” training data for fine-tuning LLMs through iterations of BOLT. Starting with the BOLT-$1138$ model, we collect top solutions from a new BO round and train two variants: 1) BOLT-$1426$, which adds all new solutions to the original BOLT-$1138$ set. 2) BOLT-$1138^{\star}$, which instead replaces an equal number of old solutions to maintain the same training set size. As shown in Table 1, both benefit from higher-quality data, suggesting “better” data significantly boosts performance. However, BOLT-$1138^{\star}$ underperforms BOLT-$1426$, which incorporates more and better data, confirming that both factors enhance model performance.
Method Best@50
LLM (BOLT-50) 87.84
LLM (BOLT-893) 82.31
LLM (BOLT-1138) 78.16
LLM (BOLT-1426) 63.68
LLM (BOLT-50) + SA 82.25
LLM (BOLT-893) + SA 63.05
LLM (BOLT-1138) + SA 61.46
LLM (BOLT-1426) + SA 61.54
@@ 5 Related Work
Language models as optimizers. Large language models (LLMs) have recently gained attention as sequence optimizers capable of tackling diverse black-box tasks where direct gradient information is unavailable or difficult to compute. LLM-based optimizers leverage the flexibility of natural language prompts to encode candidate solutions, constraints, and relevant task information. Methods like OPRO illustrates how iterative prompting can refine solutions (Yang et al., 2024; Zelikman et al., 2024), while other approaches integrate self-improving strategies that reuse high-performing LLM outputs for further fine-tuning (Shypula et al., 2024). This set of techniques has been applied to biophysical domains such as molecular design and protein engineering, where the LLM proposes mutations to enhance certain properties, as well as to program optimization tasks where the LLM speeds up code execution time (Shypula et al., 2024; Wang et al., 2024a; Madani et al., 2023).
Database optimization. Recent work has applied Bayesian optimization (BO) to improve overall database performance (Zhang et al., 2022; Nardi et al., 2019; Cereda et al., 2021) by tuning the parameters of the database configuration. As far as we are aware, Tao et al. (2025) were the first to apply BO to the specific setting of database query plan optimization considered in this paper. Other work has applied Reinforcement Learning (RL) to database query plan optimization (Marcus et al., 2019; Yang et al., 2022; Zhu et al., 2023). RL query optimizers learn from mistakes and improve performance over time. Unlike BO, however, RL requires large supervised datasets for pre-training and typically aims to minimize cumulative query latency rather than achieving the lowest possible latency.

### v2

@@ 3 Bayesian Optimization with LLM Transfer (BOLT)
Self-Augmentation.
As the fine-tuned LLM enhances few-shot generation with more optimization data, it is worth exploring whether the costly sequential BO processes can be minimized. Thus, we explore “self-improvement” methods to refine the LLM policy without the expense of additional optimization runs (Algorithm 3). Specifically, once an LLM has been fine-tuned using some of the tasks we set aside for training, we prompt it to generate additional solutions for all available tasks in that problem setting. We then score these solutions using the problem’s oracle and fine-tune the LLM again directly with this labeled self-generated data, in a manner similar to self-play in reinforcement learning or self-instruction in LLM training [11, 36]. By filtering and fine-tuning on its own best outputs, the LLM can iteratively teach itself how to propose better solutions.
@@ 4 Experiments
We evaluate BOLT on two distinct problem domains, each with a large number of related tasks. For both domains, problem definitions and solutions can be represented as strings. This allows BOLT to operate both in sequence space, where the LLM learns from optimization trajectories, and in latent space, where BO makes additional progress using LLM-sampled initializations.
@@
For the inner optimization loop, we implement a constrained version of the LOL-BO algorithm [24] using BoTorch and GPyTorch [1, 7]. Code to reproduce all results in the paper is available on GitHub: placeholder. For query optimization, we use an acquisition batch size of 1 with a budget of 4,000 oracle calls, while for peptide design, we employ a larger acquisition batch size of 50 with a budget of 200,000 oracle calls.
@@
The outer loop BOLT-$T$ models uses instruction prompting [25, 20] to guide the LLM in producing optimized sequences. Figure 4 shows the prompt used to prompt gpt-4o-mini for efficient query plans. After each optimization iteration, we augment the training set with the highest-scoring sequences from the optimization trajectory and fine-tune gpt-4o-mini on this expanded dataset. When fine-tuning the LLM for the query plan optimization task, we use OpenAI’s automatic batch size selection option. For the peptide design task, we found that using the automatic batch size option did not provide a similar boost in performance, and we use a constant batch size of $10$. For both tasks, we fine-tuned the LLM for $2$ epochs and used the default OpenAI LR multiplier hyperparameter of $1.8$. To ensure the solutions always have the correct syntax, we filter out characters that do not correspond to strings of integers or valid amino acids for the respective tasks. Figure 5 provides additional details on the fine-tuning process and prompts used for the peptide task.
Database query plan optimization.
@@
Database query plan optimization focuses on finding query plans (including join orderings and their operators) with low execution time for a given query. We take a subset of $2933$ queries from the Cardinality Estimation Benchmark introduced by Negi et al. [28], keeping $99$ queries for validation. Following Tao et al. [41], we perform BO over query plans by encoding join orders and operators as integer lists, which are then mapped to a 64-dimensional continuous latent space using the pre-trained query plan VAE from Tao et al. [41]. For the pretraining dataset, we randomly generate $1,169,890$ query plans generated based on the database schema, separated into 80/10/10 splits. For the initial “from scratch” runs with no LLM, we initialize with the set of 50 query plans used by BAO [23], an ML-powered query optimizer we use as a baseline, that produces reasonable but non-optimal plans. Subsequent runs use 50 LLM-sampled query plans per query as initialization points. All points are sampled using a temperature parameter of 0.7 unless otherwise specified. The “task description context" $C$ used to fine-tune the LLM for this task is the full SQL query string.
Antimicrobial peptide design.
@@
For the peptide design application, we are given a library of 1000 extinct, weakly antimicrobial seed peptides $S=\left\{s_{1},...,s_{L}\right\}$. A task in this setting is to take a particular seed peptide $s_{i}$ and make modifications to it to minimize the minimum inhibitory concentration (MIC) against A. Baumannii ATCC $19606$, measured in $\mu\text{mol}/L$. We created a library of $L=1000$ extinct peptides and held out the last $100$ as validation. We ensure edited peptides maintain a minimum $75\%$ similarity to the seed peptide, defined by $1-\frac{d(S,S^{\prime})}{\text{len}(S)}$, where $d$ is the Levenshtein distance between them. All of the validation peptides are at least $25\%$ different from any other peptide in the library. Although the seed peptides don’t achieve low MICs, the hope is that bacteria are less likely to have developed resistance to their variations as they come from extinct species [44]. We assess MICs with the APEX model Wan et al. [44] and utilize a VAE trained on 4.5 million amino acid sequences Torres et al. [42] to map peptides into a $256$-dimensional latent space. Initial optimization uses $1000$ randomly mutated sequences with a similarity constraint of $75\%$ to the seed. Subsequent runs utilize $1000$ LLM sampled peptides. All points are sampled using a temperature parameter of 1.0 unless otherwise specified. We use the seed amino acid sequence as the “task description context” $C$ for LLM fine-tuning.
Baselines.
@@
We compare BOLT against a range of baseline approaches. First, we compare to “from scratch,” single task LOL-BO which we will refer to as STBO, which operates without prior task knowledge. Second, we compare to a common strategy for multi task BO—see, e.g., Patacchiola et al. [30], Hakhamaneshi et al. [10], Perrone et al. [31], where a shared GP is trained on all tasks through a neural network feature extractor using the optimization trajectories from training tasks. This shared GP is then check-pointed and used on test tasks. Several papers have found success with variations of this approach. ABLR [31] uses independent Bayesian linear regression heads per task on the shared feature extractor, while FSBO [48] uses an adaptation of DKT [30]. In this paper, for the final supervised model, we use the same PPGPR model [13] as the BO inner-loop in our method as we require scalability but find this results in better performance than Bayesian linear regression or random Fourier feature models [32].
We also compare against methods that utilize an ensemble of Gaussian process experts, POGPE and SGPE [34]. POGPE utilizes an ensemble approach with one expert from each previous task, while SGPE extends this by adding an additional expert trained on data for the current task only with higher weighting. We evaluate several configurations of these methods (with 5, 10, and 20 experts) to identify optimal performance. Additionally, we evaluate two transformer-based methods: Optformer [3] and LLAMBO [19]. Optformer employs a fine-tuned transformer model (in our implementation, GPT-4o-mini) for hyperparameter optimization, while LLAMBO uses out-of-the-box LLMs (also GPT-4o-mini in our case, compared to GPT-3.5 in the original paper) for both surrogate modeling and acquisition function optimization. Due to context length limitations with these transformer-based methods, we maintain a sliding window of the last 100 oracle calls for both approaches. For LLAMBO, we impose a budget limit of 10 million input tokens per experiment to manage computational costs. Additional details on all baselines can be found in Figure 5.
@@ 4.1 Optimization results
In Figure 1, we demonstrate that initializing BO with BOLT significantly improves optimization efficiency across both domains. On the query optimization task (left), while DKT/FSBO makes improvements over STBO, the gains appear to plateau after only 20 tasks. In contrast, BOLT successfully scales to over 1400 tasks and converges to higher quality solutions faster. On the peptide design task (right), BOLT shows similarly strong performance, while DKT/FSBO struggles to take advantage of the data collected for separate templates.
The GP expert methods show mixed results. For the database task, POGPE-5 and SGPE-5 demonstrate better performance than their variations with more experts, while for the peptide task, POGPE-10 and SGPE-10 yield the best results compared to 5/20 experts. Consistent with findings from Schilling et al. [34], POGPE generally outperforms SGPE across both domains. However, both ensemble approaches are consistently outperformed by BOLT, and even fall behind STBO and DKT/FSBO in several cases. We did not run POGPE or SGPE with a larger numbers of experts as both methods scale poorly, requiring updating of a number of GPs proportional to the number of tasks.
The transformer-based methods demonstrate notable limitations in Figures 2, LABEL: and 1. Optformer achieves performance slightly worse than STBO on the database task while showing significantly poorer results on the peptide task. LLAMBO performs substantially worse across both domains, showing minimal progress during optimization. Due to its computational demands—requiring LLM inference for both surrogate modeling and acquisition—LLAMBO completed fewer than 100 optimization steps within our token budget constraints.
We find that on both tasks once BOLT reaches a sufficient scale, it begins to generate initialization data for BO that is significantly better performing than the final results found by all baseline methods, including STBO, DKT/FSBO, the GP expert methods, and transformer-based approaches.
BOLT as a one- and few-shot optimizer.
In Figure 2, BOLT demonstrates strong few-shot generalization capabilities, even achieving single-shot performance competitive with traditional approaches. In query optimization, all BOLT variants outperform the top BAO solution within 5 samples. Notably, BOLT-1138 and BOLT-1426 surpass PostgreSQL in a single sample, indicating their potential for rapid deployment in low latency scenarios. The performance of BOLT consistently improves with more iterations across both tasks, except at 50 samples on the query optimization task, where BOLT-1138 slightly outperforms BOLT-1426. This may be due to variances in LLM sample generation or training. Overall, results confirm BOLT’s robustness when scaling to thousands of tasks. We further compare our few-shot performance (Figure 2, before x-axis break) to full BO runs (Figure 2 after x-axis break). In both tasks, BOLT achieves few shot results comparable to the full BO runs.
Figure 2: Evaluating BOLT in the few shot setting and comparing to full optimization runs in both problem settings (Left: query plan optimization; Right: peptide design). In each plot, we show objective values accumulated across all validation tasks for various methods. Scatter points illustrate the few-shot performance of BOLT using different number of tasks, and relevant domain baselines (e.g., PostgreSQL, BAO for query optimization). Horizontal dashed lines indicate the performance of various full BO runs and other optimizers, shown for comparison. These results demonstrate that BOLT’s few-shot performance is often comparable to or surpasses that of full BO runs.
@@ 4.2 Ablation Studies
Method Best@50
LLM (BOLT-50) no SA 87.84
LLM (BOLT-893) no SA 82.31
LLM (BOLT-1138) no SA 78.16
LLM (BOLT-1426) no SA 63.68
LLM (BOLT-50) 82.25
LLM (BOLT-893) 63.05
LLM (BOLT-1138) 61.46
LLM (BOLT-1426) 61.54
LLM self-augmentation.
We investigate whether self-augmentation as outlined in Algorithm 3 can improve LLM performance while avoiding the computational expense of the inner-loop BO on the query plan optimization task. We apply the self-augmentation process to BOLT-1138 and BOLT-1426, generating 10 samples from each across all $2,933$ training tasks, keeping only queries that outperform the best query plan from BAO’s solutions. We then use these two dataset as additional fine-tuning to create self-augmented versions of the LLMs.
Table 1 shows that this self-augmentation yields substantial improvements even without additional tasks optimized by BO. Both self-augmented models converge to a similar performance level, achieving a summed runtime of about 62 seconds across the 100 validation queries. This convergence suggests a natural performance plateau after training on either 1,500 tasks (BOLT-1426) or 1,100 tasks plus self-generated samples (BOLT-1138+SA). The consistency of this plateau across different training approaches further demonstrates BOLT’s robustness when scaling to large task sets. This self-augmentation experiment indicates that once the LLM has been fine-tuned to sufficient performance, it can generate additional fine-tuning data, reducing the number of BO runs required. Additionally, our framework scales well with more training tasks without performance loss.
Greedy decoding.
In all other experiments, we use a constant sampling temperature of 0.7 for response generation. In Figure 3, we demonstrate that greedy decoding (i.e., setting the temperature to 0) improves one-shot performance. Results in Figure 3 show that this strategy generates a larger set of queries that surpasses the best among 50 BAO-generated queries by a wider margin.
@@
Table 2: Comparing LLMs fine-tuned with (Left) data from 1138 tasks, (Right) data from 1426 tasks, and (Middle) data from 1138 tasks, but including the extra data from BOLT-1426, and removing data from older tasks.
Impact of data quality on training.
We perform an ablation to assess the importance of using "better" versus "more" training data for fine-tuning LLMs through iterations of BOLT. Starting with the BOLT-$1138$ model, we collect top solutions from a new BO round and train two variants: 1) BOLT-$1426$, which adds all new solutions to the original BOLT-$1138$ set. 2) BOLT-$1138^{\star}$, which instead replaces an equal number of old solutions to maintain the same training set size. As shown in Table 2, both benefit from higher-quality data, suggesting “better” data boosts performance. However, BOLT-$1138^{\star}$ underperforms BOLT-$1426$, which incorporates more and better data, confirming that both factors enhance model performance.
@@ 5 Related Work
Language models as optimizers.
Large language models (LLMs) have recently gained attention as sequence optimizers capable of tackling diverse black-box tasks where direct gradient information is unavailable or difficult to compute. LLM-based optimizers leverage the flexibility of natural language prompts to encode candidate solutions, constraints, and relevant task information. Methods like OPRO illustrates how iterative prompting can refine solutions [49, 51], while other approaches integrate self-improving strategies that reuse high-performing LLM outputs for further fine-tuning [36]. This set of techniques has been applied to biophysical domains such as molecular design and protein engineering, where the LLM proposes mutations to enhance certain properties, as well as to program optimization tasks where the LLM speeds up code execution time [36, 45, 21].
Database optimization.
Recent work has applied Bayesian optimization (BO) to improve overall database performance [52, 27, 2] by tuning the parameters of the database configuration.
As far as we are aware, Tao et al. [41] were the first to apply BO to the specific setting of database query plan optimization considered in this paper. Other work has applied reinforcement learning (RL) to query plan optimization [22, 50, 53]. RL query optimizers learn from mistakes and improve performance over time. Unlike BO, however, RL requires large supervised datasets for pre-training and typically aims to minimize cumulative query latency rather than achieving the lowest possible latency.